The revenue function​ R(x) and the cost function​ C(x) for a particular product are given. These functions are valid only for the specified range of values. Find the number of units that must be produced to break even.
R(x)=200x-2x^2 C(x)= -x^2+25x+3100 0 less than or equal to x less than or equal to 100
The manufacturer must produce ______ units to break even

Answer by mananth(16946)

For break even
R(x)= C(x)
Equate and simplify.
200x-2x^2 = -x^2+25x+3100
x^2-175x +3100=0
(x-20)(x-155)=0
x= 20 or x=155
x<100
therefore x=20
The manufacturer must produce 20 units to break even